# Probing quantum complexity via universal saturation of stabilizer entropies

### v2

@@ 1 Introduction
$\alpha$-Stabilizer Rényi entropies (SREs) have been proposed to quantitatively explore nonstabilizerness harnessing their efficient numerical [Leone et al.(2022)Leone, Oliviero, and Hamma, Haug and Piroli(2023a), Lami and Collura(2023)], analytic [López and Kos(2024), Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant] and experimental accessibility [Haug and Kim(2023), Oliviero et al.(2022)Oliviero, Leone, Hamma, and Lloyd, Haug et al.(2024)Haug, Lee, and Kim, Bluvstein et al.(2022)Bluvstein, Levine, Semeghini, Wang, Ebadi, Kalinowski, Keesling, Maskara, Pichler, Greiner, et al.]. Here, the Rényi index $\alpha$ indicates the moment of the SRE. They are monotones for pure states for $\alpha\geq 2$ [Leone and Bittel(2024)], while monotonicity can be violated for $\alpha<2$ [Haug and Piroli(2023b)]. SREs relate to phases of error-corrected circuits [Niroula et al.(2023)Niroula, White, Wang, Johri, Zhu, Monroe, Noel, and Gullans], quantify the entanglement spectrum [Tirrito et al.(2024)Tirrito, Tarabunga, Lami, Chanda, Leone, Oliviero, Dalmonte, Collura, and Hamma], bound fidelity estimation [Leone et al.(2024a)Leone, Oliviero, Esposito, and Hamma, Leone et al.(2023)Leone, Oliviero, and Hamma, Hinsche et al.(2024)Hinsche, Ioannou, Jerbi, Leone, Eisert, and Carrasco], characterize the robustness of shadow tomography [Brieger et al.(2023)Brieger, Heinrich, Roth, and Kliesch] and characterize pseudorandom states [Gu et al.(2024a)Gu, Leone, Ghosh, Eisert, Yelin, and Quek, Haug et al.(2023)Haug, Bharti, and Koh]. Further, SREs characterize many-body phenomena [Gu et al.(2024b)Gu, Oliviero, and Leone] such as phase transitions [Haug and Piroli(2023a), Tarabunga et al.(2023)Tarabunga, Tirrito, Chanda, and Dalmonte], frustration [Odavić et al.(2023)Odavić, Haug, Torre, Hamma, Franchini, and Giampaolo], random matrix product states [Frau et al.(2024)Frau, Tarabunga, Collura, Dalmonte, and Tirrito, Lami et al.(2024)Lami, Haug, and De Nardis, Paviglianiti et al.(2024)Paviglianiti, Lami, Collura, and Silva, Mello et al.(2024)Mello, Santini, and Collura], localization [Turkeshi et al.(2023b)Turkeshi, Schirò, and Sierant] and out-of-time-order correlators [Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma, Garcia et al.(2023)Garcia, Bu, and Jaffe, Leone et al.(2023)Leone, Oliviero, and Hamma]. SREs also lower bound other intractable magic monotones [Haug and Piroli(2023b), Haug et al.(2024)Haug, Lee, and Kim].
@@
Our main results are summarized in Fig. 1 and the critical T-gate density and time are summarized in Table 1.
@@ 3 Clifford circuits with T-gates
where $\Theta(x)$ is the Heaviside function with $\Theta(x)=0$ for $x\leq 0$ and $\Theta(x)=1$ for $x>0$. Stabilizer states are stabilized by a commuting subgroup $G$ of $2^{N}$ Pauli strings with |$\ket{\psi}\sigma\ket{\psi}^{2}=1$ for $\sigma\in G$. The group $G$ has $N$ generators. Applying a T-gate on a stabilizer state breaks at most one generator of $G$, resulting in a state with $N-1$ generators and $2^{N-1}$ stabilizing Pauli strings. In fact, each additional T-gate can break only one generator. Thus, we find $M_{0}\leq N_{\text{T}}\ln(2)$ and $M_{0}\leq N\ln(2)$ [Leone et al.(2022)Leone, Oliviero, and Hamma]. When the T-gate is applied after a random Clifford circuit, the $T$-gate will break one of the generators of $G$ with overwhelming probability. Thus, with overwhelming probability $N_{\text{T}}\approx N$ T-gates are necessary and sufficient to reach $M_{0}=0$, thus the critical T-gate density is $q_{\text{c},0}=1$.
@@
Figure 3: Saturation of $\alpha$-SREs for random Clifford circuits doped with T-gates. a) $M_{\alpha}$ divided by $\alpha$-SRE per T-gate $M_{\alpha}^{\text{T}}$ against T-gate density $q$. The crosses denote the critical T-gate density $q_{\text{c},\alpha}$ as derived in Eq. (18). b) $M_{\alpha}$ divided by maximal SRE $M_{\alpha}^{\text{max}}$ from Eq. (17) derived for $N\gg 1$. We show $N=14$ qubits.
@@
Next, we estimate the value of the maximal SRE for $N\gg 1$. First, we recall the fact that nonstabilizerness approaches the asymptotically the maximal possible value for randomly chosen states [Liu and Winter(2022)]. Thus, by estimating the SRE of a random state for large $N$ we can approximate $M_{\alpha}^{\text{max}}$. A random state is spread out over the whole Pauli spectrum where for simplicity we assume a uniform distribution with $\bra{\psi}\sigma\ket{\psi}^{2}=2^{-N}$ for $\sigma\in\mathcal{P}/\{I\}$. While such a state is not a positive density matrix, we find from numerical studies that this spectrum is a sufficiently good approximation of an actual random Pauli spectrum for large $N$. Using this ansatz, we find
@@
This result is confirmed by the average $M_{\alpha}(\ket{\psi_{\text{Haar}}})$ for Haar random states which has been computed in Ref. [Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant], where Haar random states are known to have asymptotically maximal nonstabilizerness [Liu and Winter(2022)]. Now, in the limit $N\gg 1$ we find
@@ 4 Random basis evolution
|$\ket{\psi_{\text{c}}(t)}=U_{\text{C}}^{(0)}\prod_{k=1}^{d}(R_{z}(\frac{2}{\sqrt{d}})\otimes I_{N-1})U_{\text{C}}^{(k)}\ket{0}\,.$ (19)
Here, we can interpret $t$ as evolution time. We argue that this circuit model describes a type of (time-dependent) random evolution: In particular, if we regard one layer, it consists of transformation into a random basis with Clifford $U_{\text{C}}^{(k)}$ and $z$-rotation in the transformed basis by small angle $\theta=2t/\sqrt{d}$. This can be seen as a kind of a trotterized evolution with time-dependent Hamiltonian $H(t)$ which rapidly changes between different bases. We find numerically that the dynamics matches closely the evolution time of random Hamiltonians as shown in Appendix A.
@@
up to a time $t_{\text{c}}^{2}\approx N\ln(2)$, where it then converges to the average value of Haar random states (see Appendix B). From the bound Eq. (6) we have for $\alpha=2$ [Haug and Piroli(2023b)]
@@ 5 Random Hamiltonian evolution
Now, how does the Pauli spectrum $\beta_{\sigma}(t)^{2}$ change when the stabilizer state is evolved in time $t$? For $t=0$, the Pauli spectrum has two peaks at $\beta_{\sigma}^{2}=0$ and $\beta_{\sigma}^{2}=1$. For $t>0$, the two peaks shift and diffuse. However, we observe numerically that the two peaks remain highly concentrated even for relatively large $t$. Note that for Haar random states, the concentration has been proven [Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant]. We show the histogram of the Pauli spectrum in Fig. 5a Note that up to $t\lesssim 1$, there are two distinct peaks with a gap in between them. Note that Fig. 5a is a logarithmic plot, and the peak for small $\beta_{\sigma}^{2}$ appears broad in logarithmic space, but is actually very concentrated close to its mean value. Let us now approximate the two peaks as Delta-functions centered around their mean value. For many qubits $N\gg 1$, we can easily compute the average of each peak, i.e. $\beta_{\sigma\in G}^{2}\approx F^{2}$, and $\beta_{\sigma\notin G}^{2}\approx 2^{-N}(1-F^{2})$.
@@
Our analytic results match numerical simulations as shown in Fig. 5b for all investigated $\alpha$ and $N$. While we assumed small $t$ for the approximations, we observe that our equations match our numerical studies until the critical time where the SRE becomes maximal.
@@
Figure 5: $\alpha$-SREs of random Hamiltonian evolution. a) Pauli spectrum plotted as histogram, where we show the probability $C$ of observing the Pauli expectation values $\beta_{\sigma}^{2}=\bra{\psi}\sigma\ket{\psi}^{2}$. We show different $t$ of GUE evolution for $N=10$ and averages over 20 random instances. For clarity, we do not show the trivial identity operator $\bra{\psi}I\ket{\psi}^{2}=1$ of the spectrum. b) SRE density $m$, where dashed lines are approximations Eq. (32), Eq. (29), Eq. (30) and Eq. (31). The sudden change of the dashed line for $\alpha<1$ is due to the change from Eq. (29) to Eq. (30) at $12^{N(1-\alpha)}(1-F^{2})^{\alpha}=$ 1. c) SRE $M$ against $N$ for $t=10^{-2}$. Dashed line is fit with approximations. Our model accurately describes $\alpha$-SREs of random evolution as function of $N$ and $t$, allowing us to predict overall scaling and critical time $t_{\text{c},\alpha}$.
@@
Finally, as we show in Appendix A, the Pauli spectrum and SRE of the GUE evolution matches closely the dynamics of the random basis evolution of Sec. 4. We also observe that the SREs for both models match. Thus, we argue that the scale-invariant behavior that we shown analytically in Sec. 4 for random bases evolution also emerges for the evolution with random Hamiltonians.
@@ 6 Complexity and SRE
First, we note that for $\alpha>1$, the SRE is similar to $D_{\text{min}}$, which is the distance to the closest stabilizer state [Haug et al.(2024)Haug, Lee, and Kim]. We find that the previously proven lower bound [Haug and Piroli(2023b)] is indeed tight for random evolutions Eq. (24) which we numerically confirm in Appendix C. Note that numeric evidence shows that SREs for $\alpha>1$ also provide an $N$-independent upper bound to $D_{\text{min}}$ [Haug et al.(2024)Haug, Lee, and Kim]. As such, we argue that SREs with $\alpha>1$ probe the closeness to the nearest stabilizer state.
@@
In Appendix D, we show an explicit fidelity estimation algorithm whose efficiency is directly given by $M_{2}(\ket{\psi})$: Fist, we note that fidelity estimation works especially well when |$\ket{\psi}$ is a stabilizer state. Now, if |$\ket{\psi}$ is not a stabilizer, but close to stabilizer state, then one can use this stabilizer state as a proxy for fidelity estimation: One estimates the fidelity of $\rho$ with the closest stabilizer state $\text{argmax}_{\ket{\psi_{\text{STAB}}}}|\braket{\psi}{\psi_{\text{STAB}}}|^{2}$. We show that when |$\ket{\psi}$ is close to a stabilizer state, this is a good approximation of the fidelity, where the quality depends on $M_{(}\ket{\psi})$. In particular, we show in Appendix D that this scheme has an error $\Delta F\leq 2\sqrt{1-\exp(-M_{2}(\ket{\psi}))}$ and gives a non-trivial approximation of the fidelity as long as $M_{2}(\ket{\psi})\leq\log(4/3)$.
In contrast, SREs with $\alpha<1$ (especially $\alpha=1/2$) show behavior similar to the log-free robustness of magic LR [Howard and Campbell(2017), Rall et al.(2019)Rall, Liang, Cook, and Kretschmer, Liu and Winter(2022)] or max-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)]. They respectively relate to the negativity of the mixture of stabilizer states, or the number of superpositions of stabilizer states needed to represent a given state. LR has been used to estimate fault-tolerant state preparation complexity and relates to the complexity of Clifford based simulation algorithms. These algorithms simulate quantum circuits as Clifford circuits injected with nonstabilizer gates, where the simulation complexity commonly increase exponentially with the number of nonstabilizer gates [Bravyi et al.(2016)Bravyi, Smith, and Smolin]. In fact, $M_{1/2}$ has been used as a proxy for LR to evaluate simulation complexity of Clifford-based simulation algorithms [Rall et al.(2019)Rall, Liang, Cook, and Kretschmer]. Further, we find that the lower bound $\text{LR}\geq\frac{1}{2}M_{1/2}$ Eq. (8) is tight for random evolution (see Appendix C). Additionally, $M_{0}$ is a lower bound to the stabilizer nullity $\nu\geq M_{0}$, which characterizes the complexity of Clifford-based learning algorithms [Gu et al.(2024c)Gu, Oliviero, and Leone, Grewal et al.(2023)Grewal, Iyer, Kretschmer, and Liang, Hangleiter and Gullans(2023), Leone et al.(2024b)Leone, Oliviero, and Hamma].
@@
In contrast, SREs for random evolution shows widely different behavior depending on $\alpha$. This is because random evolution affects all Pauli strings even at short evolution times. For $\alpha>1$, $M_{\alpha>1}$ grows as $\propto t^{2}$ and converges to its maximum at $t_{\text{c},\alpha>1}\sim\sqrt{N}$. For $t=\text{const}$, we have $M_{\alpha>1}=\text{const}$ and $D_{\text{min}}=\text{const}$, i.e. the evolution is close to a stabilizer state in terms of fidelity. Thus, at $t=\text{const}$ one can efficiently certify the fidelity of random evolution (see Appendix D).
@@
Finally, we note that the complexity of tasks related to entanglement tasks can be bounded using the stabilizer nullity [Gu et al.(2024c)Gu, Oliviero, and Leone]. For small stabilizer nullity, one can efficiently do entanglement estimation, distillation and dilution. The same tasks become hard when the stabilizer nullity becomes extensive. For Cliffords doped with a constant number of T-gates we have a constant stabilizer nullity, and thus entanglement-related tasks are easy. In contrast, for random evolution the stabilizer nullity becomes already maximal for exponentially small times, where entanglement-related tasks become hard.

### v3

@@ 1 Introduction
$\alpha$-Stabilizer Rényi entropies (SREs) have been proposed to quantitatively explore nonstabilizerness harnessing their efficient numerical [Leone et al.(2022)Leone, Oliviero, and Hamma, Haug and Piroli(2023a), Lami and Collura(2023)], analytic [López and Kos(2024), Turkeshi et al.(2025)Turkeshi, Dymarsky, and Sierant] and experimental accessibility [Haug and Kim(2023), Oliviero et al.(2022)Oliviero, Leone, Hamma, and Lloyd, Haug et al.(2024)Haug, Lee, and Kim, Bluvstein et al.(2024)Bluvstein, Evered, Geim, Li, Zhou, Manovitz, Ebadi, Cain, Kalinowski, Hangleiter, et al., Haug and Tarabunga(2025)]. Here, the Rényi index $\alpha$ indicates the moment of the SRE. They are monotones for pure states for $\alpha\geq 2$ [Leone and Bittel(2024)], while monotonicity can be violated for $\alpha<2$ [Haug and Piroli(2023b)]. SREs relate to phases of error-corrected circuits [Niroula et al.(2024)Niroula, White, Wang, Johri, Zhu, Monroe, Noel, and Gullans], quantify the entanglement spectrum [Tirrito et al.(2024)Tirrito, Tarabunga, Lami, Chanda, Leone, Oliviero, Dalmonte, Collura, and Hamma], bound fidelity estimation [Leone et al.(2024a)Leone, Oliviero, Esposito, and Hamma, Leone et al.(2023)Leone, Oliviero, and Hamma, Hinsche et al.(2024)Hinsche, Ioannou, Jerbi, Leone, Eisert, and Carrasco], characterize the robustness of shadow tomography [Brieger et al.(2025)Brieger, Heinrich, Roth, and Kliesch] and characterize pseudorandom states [Gu et al.(2024a)Gu, Leone, Ghosh, Eisert, Yelin, and Quek, Haug et al.(2025)Haug, Bharti, and Koh]. Further, SREs characterize many-body phenomena [Gu et al.(2024b)Gu, Oliviero, and Leone] such as phase transitions [Haug and Piroli(2023a), Tarabunga et al.(2023)Tarabunga, Tirrito, Chanda, and Dalmonte], frustration [Odavić et al.(2023)Odavić, Haug, Torre, Hamma, Franchini, and Giampaolo], random matrix product states [Frau et al.(2024)Frau, Tarabunga, Collura, Dalmonte, and Tirrito, Lami et al.(2025)Lami, Haug, and De Nardis, Paviglianiti et al.(2024)Paviglianiti, Lami, Collura, and Silva, Mello et al.(2024)Mello, Santini, and Collura], localization [Turkeshi et al.(2023)Turkeshi, Schirò, and Sierant] and out-of-time-order correlators [Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma, Garcia et al.(2023)Garcia, Bu, and Jaffe, Leone et al.(2023)Leone, Oliviero, and Hamma]. SREs also lower bound other intractable magic monotones [Haug and Piroli(2023b), Haug et al.(2024)Haug, Lee, and Kim].
@@
Our main results are summarized in Fig. 1, while the critical T-gate density and time are shown in Table 1.
@@ 3 Clifford circuits with T-gates
where $\Theta(x)$ is the Heaviside function with $\Theta(x)=0$ for $x\leq 0$ and $\Theta(x)=1$ for $x>0$. Stabilizer states are stabilized by a commuting subgroup $G$ of $2^{N}$ Pauli strings with $\bra{\psi}\sigma\ket{\psi}^{2}=1$ for $\sigma\in G$. The group $G$ has $N$ generators. Applying a T-gate on a stabilizer state breaks at most one generator of $G$, resulting in a state with $N-1$ generators and $2^{N-1}$ stabilizing Pauli strings. In fact, each additional T-gate can break only one generator. Thus, we find $M_{0}\leq N_{\text{T}}\ln(2)$ and $M_{0}\leq N\ln(2)$ [Leone et al.(2022)Leone, Oliviero, and Hamma]. When the T-gate is applied after a random Clifford circuit, the $T$-gate will break one of the generators of $G$ with overwhelming probability. Thus, with overwhelming probability $N_{\text{T}}\approx N$ T-gates are necessary and sufficient to reach $M_{0}=0$, thus the critical T-gate density is $q_{\text{c},0}=1$.
@@
Figure 3: Saturation of $\alpha$-SREs for random Clifford circuits doped with T-gates. a) $M_{\alpha}$ divided by $\alpha$-SRE per T-gate $M_{\alpha}^{\text{T}}$ against T-gate density $q$. The crosses denote the critical T-gate density $q_{\text{c},\alpha}$ as derived in Eq. (18). b) $M_{\alpha}$ divided by maximal SRE $M_{\alpha}^{\text{max}}$ from Eq. (17) derived for $N\gg 1$. We show $N=14$ qubits and average over 20 random instances, where we note that due to self-averaging the variance in SRE is small.
@@
Next, we estimate the value of the maximal SRE for $N\gg 1$. First, we recall the fact that nonstabilizerness approaches the asymptotically the maximal possible value for randomly chosen states [Liu and Winter(2022)]. Thus, by estimating the SRE of a random state for large $N$ we can approximate $M_{\alpha}^{\text{max}}$. A random state is spread out over the whole Pauli spectrum where for simplicity we assume a uniform distribution with $\bra{\psi}\sigma\ket{\psi}^{2}=2^{-N}$ for $\sigma\in\mathcal{P}/\{I\}$. While such a state is not a positive density matrix, we find from numerical studies that this spectrum is a sufficiently good approximation of an actual random Pauli spectrum for large $N$. In fact, we find the concentration around our predicted value for highly random states as shown in Fig. 5a for $t=1.5$. Using our ansatz, we find
@@
As further confirmation, our result matches $M_{\alpha}^{\text{Haar}}$ of Haar random states up to constant corrections [Turkeshi et al.(2025)Turkeshi, Dymarsky, and Sierant], where Haar random states are known to have asymptotically maximal nonstabilizerness [Liu and Winter(2022)] (see Appendix A for explicit comparison). Now, in the limit $N\gg 1$ we find
@@ 4 Random basis evolution
|$\ket{\psi_{\text{c}}(t)}=U_{\text{C}}^{(0)}\prod_{k=1}^{d}\left(R_{z}\left(\frac{2t}{\sqrt{d}}\right)\otimes I_{N-1}\right)U_{\text{C}}^{(k)}\ket{0}\,.$ (19)
Here, we can interpret $t$ as evolution time. We argue that this circuit model describes a type of (time-dependent) random evolution: In particular, if we regard one layer, it consists of transformation into a random basis with Clifford $U_{\text{C}}^{(k)}$ and $z$-rotation in the transformed basis by small angle $\theta=2t/\sqrt{d}$. This can be seen as a kind of a trotterized evolution with time-dependent Hamiltonian $H(t)$ which rapidly changes between different bases. We find numerically that the dynamics matches closely the evolution time of random Hamiltonians as shown in Appendix B.
@@
up to a time $t_{\text{c}}^{2}\approx N\ln(2)$, where it then converges to the average value of Haar random states (see Appendix C). From the bound Eq. (6) we have for $\alpha=2$ [Haug and Piroli(2023b)]
@@ 5 Random Hamiltonian evolution
Now, how does the Pauli spectrum $\beta_{\sigma}(t)^{2}$ change when the stabilizer state is evolved in time $t$? For $t=0$, the Pauli spectrum has two peaks at $\beta_{\sigma}^{2}=0$ and $\beta_{\sigma}^{2}=1$. For $t>0$, the two peaks shift and diffuse. However, we observe numerically that the two peaks remain highly concentrated even for relatively large $t$. Note that for Haar random states, the concentration has been proven [Turkeshi et al.(2025)Turkeshi, Dymarsky, and Sierant]. We show the histogram of the Pauli spectrum in Fig. 5a Note that up to $t\lesssim 1$, there are two distinct peaks with a gap in between them. Note that Fig. 5a is a logarithmic plot, and the peak for small $\beta_{\sigma}^{2}$ appears broad in logarithmic space, but is actually very concentrated close to its mean value. Let us now approximate the two peaks as Delta-functions centered around their mean value. For many qubits $N\gg 1$, we can easily compute the average of each peak, i.e. $\beta_{\sigma\in G}^{2}\approx F^{2}$, and $\beta_{\sigma\notin G}^{2}\approx 2^{-N}(1-F^{2})$.
@@
Our analytic results match numerical simulations as shown in Fig. 5b,c for all investigated $\alpha$ and $N$. While we assumed small $t$ for the approximations, we observe that our equations match our numerical studies until the critical time where the SRE becomes maximal. We show additional numerical results and the error relative to the approximations in Appendix D.
@@
Figure 5: $\alpha$-SREs of random Hamiltonian evolution. a) Pauli spectrum plotted as histogram, where we show the probability $C$ of observing the Pauli expectation values $\beta_{\sigma}^{2}=\bra{\psi}\sigma\ket{\psi}^{2}$. We show different $t$ of GUE evolution for $N=10$ averaged over 20 random instances. For clarity, we do not show the trivial identity operator $\bra{\psi}I\ket{\psi}^{2}=1$ of the spectrum. For $t\gg 1$, Pauli spectrum is mostly concentrated around $\beta_{\sigma}^{2}=2^{-N}\approx 10^{-3}$. b) SRE $M_{\alpha}$ against $t$ for different $\alpha$ and $N=12$, where dashed lines are approximations Eq. (32), Eq. (29), Eq. (30) and Eq. (31). The sudden change of the dashed line for $\alpha<1$ is due to the change from Eq. (29) to Eq. (30) at $12^{N(1-\alpha)}(1-F^{2})^{\alpha}=$ 1. c) SRE $M_{\alpha}$ against $N$ for $t=10^{-2}$. Dashed line is fit with approximations. Our model accurately describes $\alpha$-SREs of random evolution as function of $N$ and $t$, allowing us to predict overall scaling and critical time $t_{\text{c},\alpha}$.
@@
Finally, as we show in Appendix B, the Pauli spectrum and SRE of the GUE evolution matches closely the dynamics of the random basis evolution of Sec. 4. We also observe that the SREs for both models match. Thus, we argue that the scale-invariant behavior that we shown analytically in Sec. 4 for random bases evolution also emerges for the evolution with random Hamiltonians.
@@ 6 Complexity and SRE
First, we note that for $\alpha>1$, the SRE is similar to $D_{\text{min}}$, which is the distance to the closest stabilizer state [Haug et al.(2024)Haug, Lee, and Kim]. We find that the previously proven lower bound [Haug and Piroli(2023b)] is indeed tight for random evolutions Eq. (24) which we numerically confirm in Appendix E. Note that numeric evidence shows that SREs for $\alpha>1$ also provide an $N$-independent upper bound to $D_{\text{min}}$ [Haug et al.(2024)Haug, Lee, and Kim]. As such, we argue that SREs with $\alpha>1$ probe the closeness to the nearest stabilizer state.
@@
In Appendix F, we show an explicit fidelity estimation algorithm whose efficiency is directly given by $M_{2}(\ket{\psi})$: Fist, we note that fidelity estimation works especially well when |$\ket{\psi}$ is a stabilizer state. Now, if |$\ket{\psi}$ is not a stabilizer, but close to stabilizer state, then one can use this stabilizer state as a proxy for fidelity estimation: One estimates the fidelity of $\rho$ with the closest stabilizer state $\text{argmax}_{\ket{\psi_{\text{STAB}}}}|\braket{\psi}{\psi_{\text{STAB}}}|^{2}$. We show that when |$\ket{\psi}$ is close to a stabilizer state, this is a good approximation of the fidelity, where the quality depends on $M_{2}(\ket{\psi})$. In particular, we show in Appendix F that this scheme has an error $\Delta F\leq 2\sqrt{1-\exp(-M_{2}(\ket{\psi}))}$ and gives a non-trivial approximation of the fidelity as long as $M_{2}(\ket{\psi})\leq\log(4/3)$.
In contrast, SREs with $\alpha<1$ (especially $\alpha=1/2$) show behavior similar to the log-free robustness of magic LR [Howard and Campbell(2017), Rall et al.(2019)Rall, Liang, Cook, and Kretschmer, Liu and Winter(2022)] or max-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)]. They respectively relate to the negativity of the mixture of stabilizer states, or the number of superpositions of stabilizer states needed to represent a given state. LR has been used to estimate fault-tolerant state preparation complexity and relates to the complexity of Clifford based simulation algorithms. These algorithms simulate quantum circuits as Clifford circuits injected with nonstabilizer gates, where the simulation complexity commonly increase exponentially with the number of nonstabilizer gates [Bravyi et al.(2016)Bravyi, Smith, and Smolin]. In fact, $M_{1/2}$ has been used as a proxy for LR to evaluate simulation complexity of Clifford-based simulation algorithms [Rall et al.(2019)Rall, Liang, Cook, and Kretschmer]. Further, we find that the lower bound $\text{LR}\geq\frac{1}{2}M_{1/2}$ Eq. (8) is tight for random evolution (see Appendix E). Additionally, $M_{0}$ is a lower bound to the stabilizer nullity $\nu\geq M_{0}$, which characterizes the complexity of Clifford-based learning algorithms [Gu et al.(2025)Gu, Oliviero, and Leone, Grewal et al.(2023)Grewal, Iyer, Kretschmer, and Liang, Hangleiter and Gullans(2024), Leone et al.(2024b)Leone, Oliviero, and Hamma].
@@
In contrast, SREs for random evolution shows widely different behavior depending on $\alpha$. This is because random evolution affects all Pauli strings even at short evolution times. For $\alpha>1$, $M_{\alpha>1}$ grows as $\propto t^{2}$ and converges to its maximum at $t_{\text{c},\alpha>1}\sim\sqrt{N}$. For $t=\text{const}$, we have $M_{\alpha>1}=\text{const}$ and $D_{\text{min}}=\text{const}$, i.e. the evolution is close to a stabilizer state in terms of fidelity. Thus, at $t=\text{const}$ one can efficiently certify the fidelity of random evolution (see Appendix F).
@@
Finally, we note that the complexity of tasks related to entanglement tasks can be bounded using the stabilizer nullity [Gu et al.(2025)Gu, Oliviero, and Leone]. For small stabilizer nullity, one can efficiently do entanglement estimation, distillation and dilution. The same tasks become hard when the stabilizer nullity becomes extensive. For Cliffords doped with a constant number of T-gates we have a constant stabilizer nullity, and thus entanglement-related tasks are easy. In contrast, for random evolution the stabilizer nullity becomes already maximal for exponentially small times, where entanglement-related tasks become hard.